Please help me and show step by step.
1.The polynomial of degree 4, P(x) has a root of multiplicity 2 at x=4 and roots of multiplicity 1 at x=0 and x=−3 . It goes through the point (5,32) .
Find a formula for P(x) .

Answer by rothauserc(4718)

Root x=4 of multiplicity 2 means that P(x) has a factor (x-4)^2
Roots of multiplicity 1 at x=0 means P(x) has a factor x and at
x=-3 means a factor (x+3)
****************************************************************
P(x) = x * (x-4)^2 * (x+3)
P(x) = (x^2+3x) * (x-4)^2
P(x) = (x^2+3x) * (x^2 -8x +16)
P(x) = x^4 -5x^3 -8x^2 +48x
****************************************************************
now we use point (5,32)
32 = 5^4 -5(5^3) -8(5^2) +48(5)
32 = 625 -625 -200 +240
to make this work we have to subtract 8 from the right side of =, therefore
****************************************************************
P(x) = x^4 -5x^3 -8x^2 +48x -8 and then
32 = 32